Question 1 1 pts You invest $116 today and expect to receive $163 in 17 years. What interest rate are you expecting to earn? (Put rate in 4 decimal places.)
Answer #1
Future value = Present Value * (1+r)^n
Present value is $ 116/.
Future value is $ 163/.
r is Rate of Interest = ?
n is No of years = 17
163=116*(1+r)^17
(1+r)^17= (163/116)
(1+r)^17 = 1.405172
Using trail and error, r is 2.021%
Interest rate you expect is 2.021%